OpenStudy (anonymous):

A square garden plot has an area of 24 ft2. Find the length of each side in simplest radical form. Calculate the length of each side to the nearest tenth of a foot.

jhonyy9 (jhonyy9):

so hope you know that area of a square is side time side so this mean length time length so l^2

OpenStudy (anonymous):

\[x^2=24,x=\sqrt{24}=\sqrt{4*6}=?\]

jhonyy9 (jhonyy9):

than you know l^2 = 24